OpenStudy (anonymous):

find the values of a h and k that make the equation 2x^2+8x+4= a(x-h)^2+k

OpenStudy (anonymous):

easy way. start with \[2(x^2+4x)+4\] then \[2(x+2)^2+\text{stuff}\]

OpenStudy (anonymous):

i got the "2" from half of 4. now just go to original equation and replace x by -2 since that will make (x+2)=0 and you will find k

OpenStudy (anonymous):

do you mind i have two more questions from this question

OpenStudy (anonymous):

\[2(-2)^2+8\times -2+4=8-16+4=-8+4=-4\] and that is your k

OpenStudy (anonymous):

could you help?

OpenStudy (anonymous):

so answer is \[2(x+2)^2-4\]

OpenStudy (anonymous):

sure just ask

OpenStudy (anonymous):

Matrices don't commute to work is ust a little joke, play on words.

OpenStudy (anonymous):

btw you can also "complete the square" but this is easier. @estudier i was a little slow on that one

OpenStudy (anonymous):

thanks -3x^2-12x+5= a(x-h)^2+k

OpenStudy (anonymous):

ok start with \[-3(x^2+4x)+5\] and proceed as before

OpenStudy (anonymous):

take half of 4 and write \[-3(x+2)^2+\text{stuff}\]

OpenStudy (anonymous):

to find stuff, just replace x by -2 in original expression

OpenStudy (anonymous):

since if x = -2 the terms (x+2) is 0

OpenStudy (anonymous):

this is quite confusing..

OpenStudy (anonymous):

but im sorta getting it.. :s

OpenStudy (anonymous):

\[-3(-2)^2-12\times -2+5\] \[-12+24+5\] \[12+5\] \[17\]

OpenStudy (anonymous):

hold on one sec

OpenStudy (anonymous):

suppose you have \[y=-3(x+2)^2+k\] is it not clear that if \[x=-2\] then \[y=k\]?

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

because the first term must be 0

OpenStudy (anonymous):

and don't forget these are supposed to be the same. that is you are supposed to have \[-3x^2-12x+5=-3(x+2)^2+k\]

OpenStudy (anonymous):

so if i replace x by -2 in the left hand side it is the same as replacing x by -2 in the right hand side

OpenStudy (anonymous):

on the right i will get k on the left i will get whatever number i get, so k must be that number.

OpenStudy (anonymous):

and if you do a couple the pattern will be clear. that is replace x by -2 and the first two terms are -12 and +24

OpenStudy (anonymous):

it always works this way. double the number of opposite sign

OpenStudy (anonymous):

so when i got -12+24+5 i knew it was right. in this case you get 17 so your answer is \[-3(x+2)^2+17\]

OpenStudy (anonymous):

one more... pls..